# and hierarchical hyperbolicity

### v2

@@ 2. Bounded extensions and quasi(homo)morphisms
By Proposition 2.16, the quasihomomorphism one has to extend is always such that its boundary is extendable. This shows a stark difference from the problem of extendability of bounded cohomology classes and extendability of quasimorphisms [extendable:survey, End of Section 1.3].
@@ 3. Hierarchical hyperbolicity of central extensions
    • For each $x\in X$, $g_{W}(\pi_{W}(x))=\pi_{gW}(g\cdot x)$ in $\mathcal{C}(gW)$.
@@ 3.1. Variations on the axioms
A HHG acts cofinitely on the set of domains, and two domains in the same orbit have isometric coordinate spaces. Therefore, up to enlarging $\delta$, we can and will assume that every $U\in\mathfrak{S}$ is either unbounded (meaning that $\mathcal{C}U$ is) or $\mathrm{diam}\mathcal{C}U\leq\delta$.
@@ 3.2. Action of the centre
$\operatorname{Big}\left(z\right)=\{U\in\mathfrak{S}\,|\,\mathrm{diam}\pi_{U}(\langle z\rangle\cdot x)=\infty\},$
that is, the set of coordinate spaces “witnessing” the infinite order of $z$.
It is clear that $\operatorname{Big}\left(z\right)$ does not depend on the choice of the basepoint. We have that $z$ has infinite order if and only if $\operatorname{Big}\left(z\right)\neq\emptyset$ [DHS, Proposition 6.4], and that $\operatorname{Big}\left(z\right)$ is a finite collection of pairwise orthogonal domains (this follows from combining [DHS, Lemma 6.7] with [HHS_II, Lemma 2.1]).
@@
(1) Let $\delta$ be a HHG constant for the structure. We first argue that the $K$-action on $\mathfrak{S}$ fixes every unbounded domain. Fix $x_{0}\in\mathcal{X}$ and $k\in K-\{1\}$, and let $W$ be any unbounded domain. For every $g,h\in E$,
@@ 3.3. Bounded central extensions and hierarchical hyperbolicity
One direction of Theorem 3.11 extends [uniform_undistortion, Theorem 5.14], which states that a bounded $\mathbb{Z}$-central extension of a HHG is a HHG. In turn, the argument there is a refinement of [HRSS_3manifold, Corollary 4.3], which dealt with the case with hyperbolic base and only relied on the fact that every $\mathbb{Z}$-central extension of a hyperbolic group is bounded [neumannreeves]. As we have seen in Example 2.13, this is no longer automatically true for general HHG. Even assuming acylindrical hyperbolicity of the base is not enough: indeed $H^{2}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{Z})\cong H^{2}(\mathbb{Z}^{2};\mathbb{Z})$, which produces an unbounded class. The argument is the same as Example 2.13, using for instance that the comparison map $H^{2}_{b}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{R})\to H^{2}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{R})$ is trivial [kevin, Example 4.7].
@@
$(\Leftarrow)$ Assume that the extension $E\to G$ is bounded. A finite extension of a HHG is a HHG (via the geometric action on the same space). Moreover, a bounded $\mathbb{Z}$-central extension of a HHG is a HHG, as one sees by combining [ABO, Lemma 4.15] and [uniform_undistortion, Theorem 5.14]. So by Corollary 2.10 and induction, we conclude that $E$ is a HHG.
@@
By arguing as in [ABO, Lemma 4.20], one can show that $\mathcal{C}U$ is a quasiline for every $U\in\mathcal{U}$. Indeed, $K\leq E^{\prime}$ has unbounded orbits on $\mathcal{C}U$, so $E^{\prime}$ does not act elliptically; moreover, the action is not parabolic as it is cobounded (this is because $\pi_{U}$ is coarsely surjective). By the classification of group actions on hyperbolic spaces (see e.g. [ABO, Theorem 4.2]), there must be a loxodromic element $h\in E^{\prime}$. As $K$ commutes with $h$, it must fix the endpoints $h^{+},h^{-}\in\partial\mathcal{C}U$. In particular, the action of $K$ is not parabolic, so some $z\in K$ is a loxodromic isometry of $\mathcal{C}U$ with endpoints $h^{\pm}$. Again, this means that every element in $E^{\prime}$, which commutes with $z$, must fix both $h^{+}$ and $h^{-}$; hence, as the action is cobounded, $\mathcal{C}U$ must be a quasiline.
@@
Step 3: reduction to a single big domain. In this step we replace some coordinate spaces in $\mathfrak{S}^{\prime\prime}$ to get a new HHG structure where $z$ has at most one big domain (this step is unnecessary if $z$ has finite order). Let $\operatorname{Big}\left(z\right)=\{U_{1},\ldots,U_{j}\}$, and for every $i=1,\ldots,j$ let $\phi_{i}\coloneq E^{\prime}\to\mathbb{R}$ be the Busemann quasimorphism for the action on $\mathcal{C}U_{i}$. We first claim that these quasimorphisms are linearly independent. Indeed, if, say, $\phi_{1}$ was a linear combination of the others, then there would be some function $f\colon\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ such that $\mathrm{d}_{U_{1}}(g,1)\leq f(\mathrm{d}_{U_{2}}(g,1),\ldots,\mathrm{d}_{U_{j}}(g,1))$ for every $g\in E^{\prime}$ (here we used that, as argued in Example 2.7, the absolute value of each quasimorphism coarsely coincides with the distance in the corresponding quasiline). However the elements of $\operatorname{Big}\left(z\right)$ are pairwise orthogonal, so the partial realisation axiom (10) implies that $E^{\prime}$ acts coboundedly on the product $\prod_{i=1}^{j}\mathcal{C}U_{i}$. In particular, one can find a sequence of $g_{n}\in E^{\prime}$ for which $\mathrm{d}_{U_{1}}(g_{n},1)\to\infty$ while the other projections are uniformly bounded, contradicting the existence of $f$.
@@
  • Consistency (7). Since $\{z\}$ is clean in $\mathfrak{S}^{\prime}$, no element of $\mathfrak{S}^{\prime\prime}$ is either transverse to, or nested in, any $U\in\operatorname{Big}\left(z\right)$. Hence, both consistency inequalities only involve domains from $\mathfrak{S}^{\prime\prime}-\operatorname{Big}\left(z\right)$, and therefore already held in the original structure.
@@
Before proceeding, we point out that the collection $\{z_{2},\ldots,z_{n}\}$ is still clean in the new structure. This is because we did not change the projections to any domain outside $\mathcal{U}$; hence, for every $i=2,\ldots,n$, $\operatorname{Big}\left(z_{i}\right)$ might consist of different domains but it must still be contained inside $\mathcal{U}$.
@@
We now check that $(\mathcal{Y},\overline{\mathfrak{S}})$ satisfies the axioms of an almost HHG structure.
@@
  • Finite complexity (3). By how nesting is defined, every chain $U_{1}]\sqsubsetneq\ldots\sqsubsetneq[U_{k}]$ in $\overline{\mathfrak{S}}$ lifts to a chain $U_{1}\sqsubsetneq\ldots\sqsubsetneq U_{k}$ in $\widetilde{\mathfrak{S}}$, whose length is uniformly bounded.
@@
    Moving to the second requirement, let $U],[V],[W]\in\overline{\mathfrak{S}}$ be such that $U]\sqsubsetneq[V]$ and $V]\bot[W]$. As above, we can find representatives $U\sqsubsetneq V$ and $V\bot W$, so we get that $U\bot W$ and therefore $U]\bot[W]$.
  • Containers (5). Let $V]\sqsubsetneq[W]$, and let $V^{\prime}]\sqsubseteq[V]$ be such that $\mathcal{C}V^{\prime}$ is unbounded. Let $\mathcal{Q}=\{[Q]\sqsubsetneq[W]\mid[Q]\bot[V]\}$, and notice that every element of $\mathcal{Q}$ is orthogonal to $V^{\prime}]$ as well. Fix a representative $W\in[W]$, and for every $Q]\in\mathcal{Q}$ let $Q\in[Q]$ be nested in $W$. Since $V^{\prime}$ is the unique representative of $V^{\prime}]$, it must be nested in $W$ and orthogonal to every $Q$. Thus the container axiom for $\widetilde{\mathfrak{S}}$ produces a domain $T\sqsubsetneq W$ which contains every $Q$. Hence $T]\sqsubsetneq[W]$ and contain all $Q]\in\mathcal{Q}$, so it is a container for $[V]$ inside $[W]$.
@@
[…]
@@
[…]
@@
Step 5: conclusion As argued in the Interlude above, to conclude the proof of Theorem 3.14 we are left to show that the collection $\{[z_{2}],\ldots,[z_{n}]\}\subset E^{\prime}/K$ is clean in $\overline{\mathfrak{S}}$. By how orthogonality and nesting are defined in $\overline{\mathfrak{S}}$, it is in turn enough to prove that $\operatorname{Big}\left([z_{i}]\right)\subseteq[\mathcal{U}]$, where $\mathcal{U}]=\{[U]\}_{U\in\mathcal{U}-\{U_{z}\}}$. Indeed, if $W]\in\overline{\mathfrak{S}}$ has unbounded coordinate space, and $\mathrm{diam}_{[W]}(\langle[z_{i}]\rangle\cdot[x_{0}])=\infty$ for some $x_{0}\in\mathcal{X}$, then $\mathrm{diam}_{W}(\langle z_{i}\rangle\cdot x_{0})=\infty$ as well, since the projection $q_{W}\colon\mathcal{C}W\to\mathcal{C}[W]$ is a quasi-isometry. Hence $W\in\operatorname{Big}\left(z_{i}\right)-\{U_{z}\}$, so $W]\in[\mathcal{U}]$, as required. ∎
@@
$G$ is a quotient of the $(3,3,3)$ triangle group $H$, since the above elements are all reflections and the product of any two of them has order $3$ (geometrically, this is because the dihedral angle between any two of the planes is $\pi/3$). Also $\alpha\circ\beta\circ\gamma$ is given by $(x,y,z)\mapsto(x+3,z-3,y)$, which has infinite order (as one sees by looking at the first coordinate). However, $H$ is just-infinite, meaning that its only infinite quotient is $H$ itself, so $G\cong H$. Indeed, this follows from [McCarthy_just_infinite, Proposition 9], together with the fact that the point group is $S_{3}$, realised as an irreducible subgroup of $\mathrm{GL}_{2}(\mathbb{Z})$.
@@ 4.1. The general problem
There are various cases known in the literature. For instance, [BHMS, Theorem 7.1] provides an affirmative answer if $\mathcal{B}$ consists of conjugacy representatives of all Dehn twists; [short_HHG:II] almost completely settles the conjecture for the five-punctured sphere; and [random_quot] proves hierarchical hyperbolicity of quotients by random walks, which should be thought as the “typical” quotients (see also [hhs_asdim] for similar classes of quotients). The simplest unknown case is where $\mathcal{B}$ consists of a single Dehn twist, and even in that case one already encounters the problem of showing that a certain quotient central extension is bounded.
@@
4.2. Quotients by pseudo-Anosovs, low complexity

### v3

@@ 2.2.2. Extending quasimorphisms from a union
As pointed out in Remark 2.18, in the setting of Proposition 2.16 there is a single bounded cocycle that simultaneously extends all $\delta\chi_{i}|_{N_{i}}$. It could be possible to exploit this to find the extensions $X_{i}$ with uniformly bounded defect.
@@ 3. Hierarchical hyperbolicity of central extensions
    • For each $x\in\mathcal{X}$, $g_{W}(\pi_{W}(x))=\pi_{gW}(g\cdot x)$ in $\mathcal{C}(gW)$.
@@
Remark 3.3 (On finite generation).
Our results in Section 2 only required finite generation of the kernel, not of the quotient or of the extension. However, note that HHGs are finitely generated by definition.
@@ 3.1. Variations on the axioms
A HHG acts cofinitely on the set of domains, and two domains in the same orbit have isometric coordinate spaces. Therefore, up to enlarging $\delta$, we can and will assume that every $U\in\mathfrak{S}$ is either unbounded (meaning that $\mathrm{diam}\mathcal{C}U=\infty$) or $\mathrm{diam}\mathcal{C}U\leq\delta$.
@@ 3.2. Action of the centre
$\operatorname{Big}\left(z\right)=\{U\in\mathfrak{S}\,|\,\mathrm{diam}\pi_{U}(\langle z\rangle\cdot x)=\infty\}.$
[…]
@@
(1) Let $\delta$ be a HHG constant for the structure. We first argue that the $K$-action on $\mathfrak{S}$ fixes every unbounded domain. Fix $x_{0}\in\mathcal{X}$, $k\in K-\{1\}$, and an unbounded domain $W\in\mathfrak{S}$. For every $g,h\in E$,
@@ 3.3. Bounded central extensions and hierarchical hyperbolicity
As we shall expand upon in the proof, the fact that a bounded $\mathbb{Z}$-central extension of a HHG is a HHG was essentially proven in [uniform_undistortion, Proposition 5.14]. In turn, the argument there refined [HRSS_3manifold, Corollary 4.3], which dealt with the case with hyperbolic base and only relied on the fact that every $\mathbb{Z}$-central extension of a hyperbolic group is bounded [neumannreeves]. As we have seen in Example 2.13, this is no longer automatically true for general HHG. Even assuming acylindrical hyperbolicity of the base is not enough: indeed $H^{2}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{Z})\cong H^{2}(\mathbb{Z}^{2};\mathbb{Z})$, which produces an unbounded class. The argument is the same as Example 2.13, using for instance that the comparison map $H^{2}_{b}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{R})\to H^{2}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{R})$ is trivial [kevin, Example 4.7].
@@
$(\Leftarrow)$ Assume that the extension $E\to G$ is bounded. Since a finite extension of a HHG is a HHG, via the geometric action on the same space, we can reduce to the case where $T=\{0\}$. Moreover, by Corollary 2.10 and induction on $n$, we can also assume that $K\cong\mathbb{Z}$.
Now, since $E\to G$ is bounded, Proposition 2.9 produces a quasimorphism $\phi\colon E\to K$ which is the identity on $K$. In turn, by [ABO, Lemma 4.15] there exists a Cayley graph $L$ for $E$ such that $\phi\colon L\to\mathbb{R}$ is a quasi-isometry. In particular $L$ is a quasiline on which $K$ acts loxodromically. Then the collection $\{E,G,L\}$ satisfies the assumptions of [uniform_undistortion, Proposition 5.14], which equips $E$ with a HHG structure.
@@
By arguing as in [ABO, Lemma 4.20], one can show that $\mathcal{C}U$ is a quasiline for every $U\in\mathcal{U}$. Indeed, $K\leq E^{\prime}$ has unbounded orbits on $\mathcal{C}U$, so $E^{\prime}$ does not act elliptically; moreover, the action is not parabolic as it is cobounded (this is because $\pi_{U}$ is coarsely surjective). By the classification of group actions on hyperbolic spaces (which dates back to Gromov [gromov], see e.g. [ABO, Theorem 4.2] for a more modern statement), there must be a loxodromic element $h\in E^{\prime}$. As $K$ commutes with $h$, it must fix the endpoints $h^{+},h^{-}\in\partial\mathcal{C}U$. In particular, the action of $K$ is not parabolic, so some $z\in K$ is a loxodromic isometry of $\mathcal{C}U$ with endpoints $h^{\pm}$. Again, this means that every element in $E^{\prime}$, which commutes with $z$, must fix both $h^{+}$ and $h^{-}$; hence, as the action is cobounded, $\mathcal{C}U$ must be a quasiline.
@@
Step 3: reduction to a single big domain. In this step we replace some coordinate spaces, while keeping the domain set $\mathfrak{S}^{\prime\prime}$, to get a new HHG structure where $z$ has at most one big domain (this step is unnecessary if $z$ has finite order). Let $\operatorname{Big}\left(z\right)=\{U_{1},\ldots,U_{j}\}$, and for every $i=1,\ldots,j$ let $\phi_{i}\coloneq E^{\prime}\to\mathbb{R}$ be the Busemann quasimorphism for the action on $\mathcal{C}U_{i}$. We first claim that these quasimorphisms are linearly independent. Indeed, if, say, $\phi_{1}$ was a linear combination of the others, then there would be some function $f\colon\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ such that $\mathrm{d}_{U_{1}}(g,1)\leq f(\mathrm{d}_{U_{2}}(g,1),\ldots,\mathrm{d}_{U_{j}}(g,1))$ for every $g\in E^{\prime}$ (here we used that, as argued in Example 2.7, the absolute value of each quasimorphism coarsely coincides with the distance in the corresponding quasiline). However the elements of $\operatorname{Big}\left(z\right)$ are pairwise orthogonal, so the partial realisation axiom (10) implies that $E^{\prime}$ acts coboundedly on the product $\prod_{i=1}^{j}\mathcal{C}U_{i}$. In particular, one can find a sequence of $g_{n}\in E^{\prime}$ for which $\mathrm{d}_{U_{1}}(g_{n},1)\to\infty$ while the other projections are uniformly bounded, contradicting the existence of $f$.
@@
  • Consistency (7). Since $\{z\}$ is clean in $\mathfrak{S}^{\prime\prime}$, no element of $\mathfrak{S}^{\prime\prime}$ is either transverse to, or nested in, any $U\in\operatorname{Big}\left(z\right)$. Hence, both consistency inequalities only involve domains from $\mathfrak{S}^{\prime\prime}-\operatorname{Big}\left(z\right)$, and therefore already held in the original structure.
@@
Before proceeding, we point out that the collection $\{z_{2},\ldots,z_{n}\}$ is still clean in the new structure. This is because we did not change the projections to any domain outside $\mathcal{U}$; hence, for every $i=2,\ldots,n$, $\operatorname{Big}\left(z_{i}\right)$ must still be contained inside $\mathcal{U}$ (though it might differ from what it was before Step 3).
@@
We now check that $(\mathcal{Y},\overline{\mathfrak{S}})$ is a HHG structure for $E^{\prime}/Z$.
@@
  • Finite complexity (3). By how nesting is defined, every chain $U_{1}]\sqsubsetneq\ldots\sqsubsetneq[U_{k}]$ in $\overline{\mathfrak{S}}$ lifts to a chain $U_{1}\sqsubsetneq\ldots\sqsubsetneq U_{k}$ in $\widetilde{\mathfrak{S}}\subseteq\mathfrak{S}^{\prime\prime}$, and the finite complexity axiom for $(\mathcal{X},\mathfrak{S}^{\prime\prime})$ uniformly bounds the length of the latter chain.
@@
    Regarding the second requirement, let $U],[V],[W]\in\overline{\mathfrak{S}}$ be such that $U]\sqsubsetneq[V]$ and $V]\bot[W]$. As above, we can find representatives $U\sqsubsetneq V$ and $V\bot W$, so we get that $U\bot W$ and therefore $U]\bot[W]$.
  • Containers (5). Let $V]\sqsubsetneq[W]$, and let $V^{\prime}]\sqsubseteq[V]$ be such that $\mathcal{C}V^{\prime}$ is unbounded. Let $\mathcal{Q}=\{[Q]\sqsubsetneq[W]\mid[Q]\bot[V]\}$, and notice that every element of $\mathcal{Q}$ is orthogonal to $V^{\prime}]$ as well. Fix a representative $W\in[W]$, and for every $Q]\in\mathcal{Q}$ let $Q\in[Q]$ be nested in $W$. Since $V^{\prime}$ is the unique representative of $V^{\prime}]$, it must be nested in $W$ and orthogonal to every $Q$. Thus the container axiom for $\mathfrak{S}^{\prime\prime}$ produces a domain $T\sqsubsetneq W$ which contains every $Q$. Hence $T]\sqsubsetneq[W]$ and contain all $Q]\in\mathcal{Q}$, so it is a container for $[V]$ inside $[W]$.
@@
  • Consistency, bounded geodesic image, and large links (7)-(9) Since $q_{W}$ is a uniform quasi-isometry for every $W\in\overline{\mathfrak{S}}$, all three axioms hold as they were true in $(\mathcal{X},\mathfrak{S}^{\prime\prime})$, plus the fact that we defined (relative) projections via the original (relative) projections.
@@
  • Uniqueness (11). Let $x],[y]\in\mathcal{Y}$, and let $r>0$ be a constant such that $\mathrm{d}_{[W]}([x],[y])\leq r$ for all $W]\in\overline{\mathfrak{S}}$. Then any $x\in[x]$ and $y\in[y]$ have uniformly close projections to all $W\in\widetilde{\mathfrak{S}}$ (again, because every $q_{W}$ is a uniform quasi-isometry). Furthermore, since $Z$ acts coboundedly on $\mathcal{C}U_{z}$, we can replace $y$ by a $K$-translate and uniformly bound $\mathrm{d}_{U_{z}}(x,y)$ as well; notice that this replacement preserves the distance between $x$ and $y$ to every $W\in\widetilde{\mathfrak{S}}$ up to a bounded error, since $Z$-orbits have uniformly bounded projections to $\mathcal{C}W$. Then the uniqueness axiom for $(\mathcal{X},\mathfrak{S}^{\prime\prime})$ yields that $x$ and $y$ are uniformly close in $\mathcal{X}$, and therefore $\mathrm{d}_{\mathcal{Y}}([x],[y])\leq\mathrm{d}_{\mathcal{X}}(x,y)$ is uniformly bounded.
@@
Step 5: conclusion As argued in the Interlude above, to conclude the proof of Theorem 3.15 we are left to show that the collection $\{[z_{2}],\ldots,[z_{n}]\}\subset E^{\prime}/K$ is clean in $\overline{\mathfrak{S}}$. By how orthogonality and nesting are defined in $\overline{\mathfrak{S}}$, it is in turn enough to prove that $\operatorname{Big}\left([z_{i}]\right)\subseteq[\mathcal{U}]$, where $\mathcal{U}]=\{[U]\}_{U\in\mathcal{U}-\{U_{z}\}}$. Indeed, if $W]\in\overline{\mathfrak{S}}$ is unbounded, and $\mathrm{diam}_{[W]}(\langle[z_{i}]\rangle\cdot[x_{0}])=\infty$ for some $x_{0}\in\mathcal{X}$, then $\mathrm{diam}_{W}(\langle z_{i}\rangle\cdot x_{0})=\infty$ as well, since the projection $q_{W}\colon\mathcal{C}W\to\mathcal{C}[W]$ is a quasi-isometry. Hence $W\in\operatorname{Big}\left(z_{i}\right)-\{U_{z}\}$, so $W]\in[\mathcal{U}]$, as required. ∎
@@
$G$ is a quotient of the $(3,3,3)$ triangle group $H$, since the above elements are all reflections and the product of any two of them has order $3$ (geometrically, this is because the dihedral angle between any two of the planes is $\pi/3$). Also $\alpha\circ\beta\circ\gamma$ is given by $(x,y,z)\mapsto(x+3,z-3,y)$, which has infinite order (as one sees by looking at the first coordinate). However, $H$ is just-infinite, meaning that whenever $H$ surjects onto an infinite group $G$ then $H\cong G$. Indeed, this follows from [McCarthy_just_infinite, Proposition 9], together with the fact that the point group is $S_{3}$, realised as an irreducible subgroup of $\mathrm{GL}_{2}(\mathbb{Z})$. This shows that $G$ is isomorphic to the $(3,3,3)$ triangle group.
@@ 4.1. The general problem
There are various cases known in the literature. For instance, [BHMS, Theorem 7.1] provides an affirmative answer if $\mathcal{B}$ consists of conjugacy representatives of all Dehn twists; [short_HHG:II] almost completely settles the conjecture for the five-punctured sphere; and [random_quot] proves hierarchical hyperbolicity of quotients by random walks, which should be thought as the “typical” quotients (see also [hhs_asdim] for similar classes of quotients, including some of those considered in [claymangahas]). The simplest unknown case is where $\mathcal{B}$ consists of a single Dehn twist, and even in that case one already encounters the problem of showing that a certain quotient central extension is bounded.
@@
4.2. Quotients by pseudo-Anosovs